OpenStudy (anonymous):

A meat market combines meat that sells for $6.50/lb with 75 pounds of meat that sells for $1.50/lb to create a mixture tht sells for $2.50/lb. Part A: Write an equation to determine the amount of $6.50/lb meat in the meat markets mixture

OpenStudy (pfenn1):

The cost of any mixture can be found by adding the cost of one ingredient with the cost of the second ingredient. The cost of any ingredient can be found by multiplying the amount (lbs) of ingredient multiplied by its unit cost ($/lb).

OpenStudy (anonymous):

soo 6.50+1.50*75lb*2.20

OpenStudy (pfenn1):

So if x is the number of pounds of meat that sells for $6.50/lb, can you write an equation that calculated the total cost of the resultant mixture? Remember....you are trying to calculate the number of lbs of meat that sells for $6.50/lb needed to make a mixture that sells for $2.50/lb. The cost equals the unit price ($/lb) multiplied by the amount of meat(lb). What would be the cost of the 75 lbs of meat that sells for $1.50/lb?

OpenStudy (anonymous):

idk

OpenStudy (pfenn1):

The cost of the 75 lbs of $1.50/lb meat = 75(1.50). Likewise the cost of the $6.50/lb meat = x(6.50) where x is the amount (lb) of $6.50/lb meat. So what is the expression for the cost of the $2.50/lb meat?

OpenStudy (anonymous):

C

OpenStudy (anonymous):

Idk whut is it lol im horible at this

OpenStudy (pfenn1):

The expression for the cost of the $2.50/lb meat = the price of the $2.50/lb meat (2.50) multiplied by the total amount of the $2.50/lb meat (75+x).

OpenStudy (anonymous):

soo 2.50/lb 75 lb = 2.50 (2.50) (75+x)

OpenStudy (pfenn1):

So the entire equation would be that the cost of the $6.50/lb meat plus the cost of the $1.50/lb meat should equal the cost of the $2.50/lb meat.\[6.50x+1.50(75)=2.50(x+75)\]In other words, if you mix x lbs of $6.50/lb meat with 75 lbs of $1.50/lb meat, you will end up with (x+75) lbs of $2.50/lb meat. Does that make any sense to you?

OpenStudy (anonymous):

yea so thts the equation

OpenStudy (anonymous):

Part b- how many pounds of the 6.50/lb meat are used in the mixture?

OpenStudy (pfenn1):

To solve for how many pounds of the 6.50/lb meat are used, you have to solve the equation for x.\[6.50x+1.50(75)=2.50(x+75)\]\[6.50x+112.5=2.50x+187.5\]\[(6.50-2.50)x=187.5-112.5\]

OpenStudy (pfenn1):

Do you know how to continue to solve this?

OpenStudy (anonymous):

(6.50-2.50)x = 4x 187.5-112.5= 75

OpenStudy (anonymous):

X= $18.75

OpenStudy (pfenn1):

Great except x in given in units of lbs, not dollars. In other words if you mix 18.75 lbs of $6.50/lb meat with 75 lbs of $1.50/lb meat you will end up with 93.75 lbs of $2.50/lb meat. Okay?

OpenStudy (anonymous):

k